OpenStudy (anonymous):

What is the point-slope (0,6) (1,3)

OpenStudy (anonymous):

@alekos

OpenStudy (anonymous):

@robtobey @radar

OpenStudy (anonymous):

to get slope subtract 6-3 and 0-1 then do 3/-1=-3/1 so slope is -3

OpenStudy (anonymous):

it says the y itercept is 0,6 sopoint slope=-3x+6

OpenStudy (anonymous):

Don't forget to medal and fan if you already haven't. Plus send me a message if you need more help @lorenahue43

OpenStudy (anonymous):

no point slope as in : y-y1=m(x-x1)

OpenStudy (anonymous):

I will if you put it in point slope:)

OpenStudy (anonymous):

point slope is y=mx+b

OpenStudy (anonymous):

are there choices

OpenStudy (anonymous):

yes .

OpenStudy (anonymous):

can i see them

OpenStudy (anonymous):

(y-3)=-3(x-1) (y+3)=3(x+1) (y+3)=-3(x+1) (y-3)=3(x+6)

OpenStudy (anonymous):

it would be A then

OpenStudy (anonymous):

if i open anoth thread will you answer another for a medal

OpenStudy (anonymous):

sure give me 5 mins

OpenStudy (anonymous):

k